OpenStudy (anonymous):

evaluate the polynomial for x=-7. 2x^2-3x+9

OpenStudy (anonymous):

Just plug in -7 as x!

OpenStudy (zehanz):

So you have to calculate: 2*(-7)²-3*-7+9. What have you got?

OpenStudy (anonymous):

i got -110

OpenStudy (anonymous):

I got 128

OpenStudy (anonymous):

just done by head dont trust me

OpenStudy (zehanz):

2*(-7)²=2*49=98 -3*-7=21 So 98+21+9=128

OpenStudy (mertsj):

\[2(-7)^2-3(-7)+9=2(49)+21+9=98+21+9=128\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Always remember WHATEVER squared is always positive

OpenStudy (anonymous):

also - · - = +

OpenStudy (mertsj):

yw